if f(x)=f(x+1/x+2),then f(-x)=f(-(x+1)/(x+2)) or f(1-x/2-x)

Grade:12

1 Answers

Askiitians Expert Ankit Jain- IIT Bombay
13 years ago

Hello Vishu,

 

Nice question you asked. I think many people get confused over the same.

In general the rule is

 

IF  f(x) = g(x)  then

f(alpha*x) = g(alpha*x).     Note in this problem  g(x) = f(x+1/x+2)  and alpha =-1.

So the correct answer is f(1-x/2-x). You just x to alpha*x , you don't apply any brackets etc etc...

 

Hope this answers your query.

Please feel free to post as many doubts on our disucssion forum as you can.  We are all IITians and here to help you in your IIT JEE preparation.

All the best!!


Regards

Askiitians Expert

Ankit Jain- IIT Bombay